# Reinforcement Learning from Bagged Reward

### v2

@@ 4.1 Reward Redistribution based on Bidirectional Attention Mechanism
Building on the foundation of Theorem 1 that optimal policies in the BRMDP and MDP are equivalent, our focus shifts to the crucial process of redistributing rewards within a bag from the BRMDP. To capture the contextual influence of each instance within the sequence, the Causal Transformer Vaswani et al. (2017) is a natural choice as a sequential prediction model. Traditionally, Transformers in RL are used in a unidirectional manner Chen et al. (2021); Janner et al. (2021a); Micheli et al. (2023), where only previous instances influence the current prediction due to the unobservability of future instances. However, given that our bagged rewards are non-Markovian and both preceding and subsequent instances affect the contribution of the current instance to the bagged reward, a bidirectional attention mechanism Seo et al. (2016); Vaswani et al. (2017); Devlin et al. (2019) becomes pivotal. This mechanism connects both past and future instances within a bag, enabling a more comprehensive understanding of contextual influences. By quantitatively evaluating the contribution of each instance, the bidirectional attention mechanism facilitates nuanced and effective reward redistribution.
4.2 Proposed Approach
@@
We introduce the Reward Bag Transformer (RBT), a novel approach designed for the BRMDP framework. The RBT is engineered to comprehend the complex dynamics of the environment through bags and to precisely predict instance-level rewards, facilitating effective reward redistribution.
@@
Causal Transformer for Reward Representation. Referring to Fig. 3, the RBT comprises a Causal Transformer (Vaswani et al., 2017; Radford et al., 2018), which maintains the chronological order of state-action pairs Chen et al. (2021); Janner et al. (2021a). For each time step $t$ in a sequence of $M$ time steps, the Causal Transformer, represented as a function $f$, processes the input sequence $\sigma=\{s_{0},a_{0},\ldots,s_{M-1},a_{M-1}\}$, generating the output $\{x_{t}\}_{t=0}^{M-1}=f(\sigma)$. By aligning the output head $x_{t}$ with the action token $a_{t}$, we directly model the consequence of actions, which are pivotal in computing immediate rewards and predicting subsequent states, which in turn aids the model in better understanding the environmental dynamics.
@@
Bidirectional Attention Layer for Reward Redistribution. Once we have obtained the output embeddings $\{x_{t}\}_{t=0}^{M-1}$, for reward prediction, they pass through a bidirectional attention layer to produce $\{\hat{r}_{t}\}_{t=0}^{M-1}$, where $\hat{r}_{t}\equiv\hat{r}_{\theta}(s_{t},a_{t})$ with $\theta$ being the RBT parameters. This layer addresses the unidirectional limitation of the Causal Transformer architecture Vaswani et al. (2017); Radford et al. (2018), integrating past and future contexts for enhanced reward prediction accuracy. For state prediction, $x_{t}$ is input into a state linear decoder, yielding the predicted next state $\hat{s}_{t+1}\equiv\hat{s}_{\theta}(s_{t},a_{t})$.
@@
The core of the RBT architecture is its bidirectional attention mechanism. For each output embedding $x_{t}$, we apply three different linear transformations to obtain embeddings for query $\mathbf{q}_{t}\in\mathbb{R}^{d}$, key $\mathbf{k}_{t}\in\mathbb{R}^{d}$, and value $v_{t}\in\mathbb{R}$. Then the instance-level reward is calculated by
@@
$\hat{r}_{t}=\sum_{\ell=0}^{M-1}\texttt{softmax}(\frac{\{\langle\mathbf{q}_{t},\mathbf{k}_{t^{\prime}}\rangle\}_{t^{\prime}=0}^{M-1}}{\sqrt{d}})_{\ell}\cdot v_{\ell},$ (4)
@@
where $d$ is the embedding dimension of the key. The rescaling operation is used to prevent extremely small gradients as in Vaswani et al. (2017). This mechanism enables the RBT to consider both the immediate and contextual relevance of each state-action pair in the trajectory when predicting rewards.
@@
4:     Sample batches from $\mathcal{D}$ and estimate bag loss based on Eq. (7).
5:     Update RBT parameters $\theta$ based on the estimated loss.
6:     Relabel rewards in $\mathcal{D}$ using the updated RBT.
7:     Optimize policy using the relabeled data by off-the-shelf RL algorithms (such as SAC).
8:  end for
@@ 4.3 Learning Objectives
Reward Prediction. The RBT is trained to ensure that, for each reward bag, the sum of predicted instance-level rewards matches the total bagged reward. This is vital for maintaining the integrity of the reward structure in the BRMDP framework. The loss function for this objective is expressed as
$\mathcal{L}_{\mathrm{r}}(\theta;B_{i,n_{i}})=\left(\sum_{t=i}^{i+n_{i}-1}\hat{r}_{t}-R(B_{i,n_{i}})\right)^{2},$ (5)
where $B_{i,n_{i}}$ represents the reward bag starting at time step $i$ with length $n_{i}$, and $R(B_{i,n_{i}})$ is the total reward for this bag. This formulation encourages the RBT to learn a nuanced distribution of rewards across states and actions within a bag. At the same time, it ensures that the sum of redistributed rewards matches the total bagged reward, maintaining consistency as per Theorem 1.
State Transition Forecasting. Alongside reward prediction, the RBT is tasked with accurately predicting the next state in the environment given the current state and action. This capability is crucial for understanding the dynamics of the environment. The corresponding loss function is:
@@
Composite Loss. The final learning objective combines the reward and state prediction losses:
@@
The RBT’s dual predictive capacity is its key advantage, enabling precise reward redistribution to individual instances and forecasting the next state. This leverages environmental dynamics for enhanced reward distribution as experimentally evidenced in the ablation study in Appendix D.2. Integrated with off-the-shelf RL algorithms such as SAC Haarnoja et al. (2018), the RBT can enhance policy learning within the BRMDP framework, as outlined in Algorithm 1.
@@ 5 Experiment
In the following experimental section, we scrutinize the efficacy of our proposed method using benchmark tasks from both the MuJoCo Brockman et al. (2016) and the DeepMind Control Suite Tassa et al. (2018) environments, focusing on scenarios with bagged rewards. We assess the performance of our method to understand its overall effectiveness and examine whether the RBT reward model accurately predicts rewards.
5.1 Compare with SOTA Methods
Experiment Setting. We evaluated our method on eight benchmark tasks from both the MuJoCo locomotion suite, including Ant-v2, Hopper-v2, HalfCheetah-v2, and Walker2d-v2, and the DeepMind Control Suite, including cheetah-run, quadruped-walk, fish-upright, and cartpole-swingup. Differing from standard environments where rewards are assigned at each step, our approach involved assigning a cumulative reward at the end of each bag while assigning a reward of zero to all other state-action pairs within the bag. The maximum length for each episode was fixed at 1000 steps across all tasks.
Baselines. In the comparative analysis, our framework was rigorously evaluated against several leading algorithms in the domain of RL with delayed reward:
@@ 5.1 Compare with SOTA Methods
  • IRCR Gangwani et al. (2020): It adopted a non-parametric uniform reward redistribution approach by using the sum of immediate per-step rewards as a stand-in for trajectory returns. We have adapted IRCR for bagged reward setting.
@@
  • HC Han et al. (2022): The HC-decomposition framework was utilized to train the policy using a value function that operates on sequences of data. This approach decoupled the value function approximation task for the current step from the historical trajectory. We employed the code as provided by the original paper.
While methods like RUDDER Arjona-Medina et al. (2019) and Align-RUDDER Patil et al. (2022) are known for addressing the problem of trajectory feedback, previous studies Gangwani et al. (2020); Ren et al. (2021); Zhang et al. (2023b) have shown superior performance using referenced methods. Additionally, since Align-RUDDER relies on successful trajectories for scoring state-action pairs, which is impractical in MuJoCo Patil et al. (2022), we ultimately excluded both methods from our comparison. Besides, detailed descriptions of the model parameters and hyper-parameters used during training are provided in Appendix C. More experimental results are included in Appendix D.1.
Figure 4: Performance comparison in MuJoCo (top row) and DeepMind Control Suite (bottom row) environments with six different fixed-length reward bag settings (5, 25, 50, 100, 200, and 500) and trajectory feedback (labeled as $9$). The mean and standard deviation are displayed over 6 trials with different random seeds across a total of 1e6 time steps.
@@
5.1.1 Fixed-Length Reward Bags
In the fixed-length reward bag experiment, we conducted experiments with six bag lengths (5, 25, 50, 100, 200, and 500) and trajectory feedback (labeled as $9$) across each environment. The experimental design followed the problem setting, which assumed that the conclusion of one reward bag directly preceded the beginning of the next. It aimed to illustrate the influence of varying bag lengths on the results, providing insight into how bag size affected the performance of the learning algorithm within these environments.
@@ 5.1.1 Fixed-Length Reward Bags
As shown in Fig. 4, the SAC method, using bagged rewards directly from the environment, suffers from a lack of guidance in agent training due to missing reward information. This issue worsens with longer bag lengths, indicating that increased reward sparsity leads to less effective policy optimization. The IRCR and RRD methods, treating rewards uniformly within a reward bag, outperform SAC, suggesting benefits from even approximate guidance. However, notable variance in their results indicates potential consistency and reliability issues. The LIRPG exhibits subdued performance across tasks, as it is proposed to solve sparse reward problems (Zheng et al., 2018), which is Markovian and does not align the reward bag setting. The HC method excels only with shorter bag lengths, suggesting that this value function modification method struggles to utilize information from longer sequences. The proposed RBT method consistently outperforms the other approaches across all the environments and bag lengths, showing that it is not only well-suited for environments with short reward bags but also capable of handling large reward bag scenarios. This demonstrated the capability of RBT to learn from the sequence of instances and, by integrating bagged reward information, accurately allocate rewards to instances, thereby guiding better policy training.
5.1.2 Arbitrary Reward Bags
@@
To validate the effectiveness of our approach under more complex conditions, we designed an experiment that allowed for overlaps or gaps between reward bags, and the length of each bag was no longer fixed. This setup simulated more realistic scenarios and tested the robustness of our method.
The results, as detailed in Table 1, affirm the superior performance of the proposed RBT method in these complex reward settings. Notably, other baseline methods that rely on reward model training experience a significant drop in effectiveness, primarily due to sample scarcity impacting model accuracy and policy learning. In contrast, the IRCR method, which distributes rewards uniformly without a model, maintains its efficacy in some environments. This outcome suggests that while approximate rewards can still guide policy learning, incorrect rewards can be highly detrimental. The consistent success of our approach in various reward bag scenarios indicates its potential for application to a broader range of reward structures, highlighting its versatility and robustness in handling more intricate reward dynamics.
@@ 6 Conclusion
In this paper, we introduce a general learning framework, Learning from Bagged Rewards (RLBR), and make theoretical connections between our learning objectives and traditional MDPs to ensure the justification of our approach. Building on this problem, we propose a Transformer-based reward model, the Reward Bag Transformer (RBT), to efficiently redistribute rewards by interpreting contextual information within bags as well as understanding environmental dynamics. The efficacy of RBT is demonstrated through extensive experiments, outperforming existing delayed-reward methods in various reward bag scenarios. Besides, our case studies highlight the RBT’s ability to effectively reallocate rewards, maintaining fidelity to the original MDP structure.
While the sum-form bagged reward currently integrated into RLBR addresses most scenarios, as evidenced by previous studies on trajectory feedback (Ren et al., 2021; Arjona-Medina et al., 2019; Zheng et al., 2018), there is value in exploring other forms of reward aggregation. Future research will investigate alternative reward structures, such as maximum values or complex combinations of latent rewards, to better capture the nuances of dynamic real-world environments. This exploration aims to enhance the adaptability and effectiveness of the RLBR framework in a broader range of applications.
@@ Appendix A Broader Impact
In this work, we introduce the problem of Reinforcement Learning from Bagged Reward (RLBR), and propose the Reward Bag Transformer to address it. On the one hand, we recognize that these techniques could raise some potential issues. As collecting bagged rewards is much more convenient and natural than gathering instance-level rewards, this could lead to some risks of abusing unauthorized data. On the other hand, we believe that developing these techniques is still necessary for the importance of solving reinforcement learning tasks with bag-level feedback. Furthermore, there are many techniques for preserving data privacy, which can be compatible with our approach to avoid such problems.

### v3

@@ 4.2 Reward Redistribution based on Bidirectional Attention Mechanism
Building on Theorem 1 that optimal policies in the BRMDP and the MDP after reward redistribution are equivalent, our focus shifts to the crucial process of redistributing rewards within a bag. To capture the contextual influence of each instance within the sequence, the Causal Transformer Vaswani et al. (2017) is a natural choice as a sequential prediction model. Traditionally, Transformers in RL are used in a unidirectional manner Chen et al. (2021); Janner et al. (2021); Micheli et al. (2023), where only previous instances influence the current prediction due to the unobservability of future instances. However, since our bagged rewards are generally non-Markovian, both preceding and subsequent instances influence the contribution of the current instance to the bagged reward, understanding the relationships among instances within a bag is therefore pivotal. This insight leads us to employ a bidirectional attention mechanism Seo et al. (2016); Vaswani et al. (2017); Devlin et al. (2019) to capture these relationships. This mechanism connects both past and future instances within a bag, enabling a more comprehensive understanding of contextual influences. By quantitatively evaluating the contribution of each instance, the bidirectional attention mechanism facilitates nuanced and effective reward redistribution, as experimentally demonstrated in Section 5.
4.3 Reward Bag Transformer
@@
We introduce the Reward Bag Transformer (RBT), a novel reward redistribution method for the BRMDP. The RBT is engineered to comprehend the complex dynamics of the environment through bags and to precisely predict instance-level rewards, facilitating effective reward redistribution.
4.3.1 Causal Transformer for Reward Representation
@@
Referring to Fig. 2, the RBT comprises a Causal Transformer (Vaswani et al., 2017; Radford et al., 2018), which maintains the chronological order of state-action pairs Chen et al. (2021); Janner et al. (2021). For each time step $t$ in a sequence of $M$ time steps, the Causal Transformer, represented as a function $f$, processes the input sequence $\sigma=\{s_{0},a_{0},\ldots,s_{M-1},a_{M-1}\}$, generating the output $\{x_{t}\}_{t=0}^{M-1}=f(\sigma)$. By aligning the output head $x_{t}$ with the action token $a_{t}$, we directly model the consequence of actions, which are pivotal in computing immediate rewards and predicting subsequent states, thereby helping the model better understand environmental dynamics.
4.3.2 Bidirectional Attention Layer for Reward Redistribution
@@
Once we have obtained the output embeddings $\{x_{t}\}_{t=0}^{M-1}$, for reward prediction, they pass through a bidirectional attention layer to produce $\{\hat{r}_{t}\}_{t=0}^{M-1}$, where $\hat{r}_{t}\equiv\hat{r}_{\theta}(s_{t},a_{t})$ with $\theta$ being the RBT parameters. This layer addresses the unidirectional limitation of the Causal Transformer architecture Vaswani et al. (2017); Radford et al. (2018), integrating past and future contexts for enhanced reward prediction accuracy. For state prediction, $x_{t}$ is input into a state linear decoder, yielding the predicted next state $\hat{s}_{t+1}\equiv\hat{s}_{\theta}(s_{t},a_{t})$.
@@
The core of the RBT architecture is its bidirectional attention mechanism. For each output embedding $x_{t}$, we apply three different linear transformations to obtain embeddings for query $\mathbf{q}_{t}\in\mathbb{R}^{d}$, key $\mathbf{k}_{t}\in\mathbb{R}^{d}$, and value $v_{t}\in\mathbb{R}$, where $d$ is the embedding dimension of the key. Then the instance-level reward is calculated by
@@
$\hat{r}_{t}=\sum_{\ell=0}^{M-1}\texttt{softmax}(\frac{\{\langle\mathbf{q}_{t},\mathbf{k}_{t^{\prime}}\rangle\}_{t^{\prime}=0}^{M-1}}{\sqrt{d}})_{\ell}\cdot v_{\ell}.$ (4)
@@
The rescaling operation is used to prevent extremely small gradients as in Vaswani et al. (2017). This mechanism enables the RBT to consider both the immediate and contextual relevance of each instance in a bag when predicting rewards.
@@
4:     Sample batches from $\mathcal{D}$.
5:     Estimate bag loss based on Eq. (7).
6:     Update RBT parameters $\theta$ based on the estimated loss.
7:     Relabel rewards in $\mathcal{D}$ using the updated RBT.
8:     Optimize policy using the relabeled data by off-the-shelf RL algorithms (such as SAC Haarnoja et al. (2018)).
9:  end for
@@ 4.4 Learning Objectives
4.4.1 Reward Prediction
The RBT is trained to ensure that, for each reward bag, the sum of predicted instance-level rewards matches the total bagged reward. This is vital for maintaining the integrity of the reward structure in the BRMDP framework. The loss function is expressed as
$\mathcal{L}_{\mathrm{r}}(\theta;B_{i,n_{i}})=\left(\sum_{t=i}^{i+n_{i}-1}\hat{r}_{t}-R(B_{i,n_{i}})\right)^{2}.$ (5)
This formulation encourages the RBT to learn a nuanced distribution of rewards across states and actions within a bag. At the same time, it ensures that the sum of redistributed rewards matches the total bagged reward, maintaining consistency as per Theorem 1.
4.4.2 State Transition Forecasting
Alongside reward prediction, the RBT is tasked with accurately predicting the next state in the environment given the current state and action. This capability is crucial for understanding the dynamics of the environment. The corresponding loss function is:
@@ 4.4.2 State Transition Forecasting
4.4.3 Composite Loss
The final learning objective combines the reward and state prediction losses:
@@ 4.4.3 Composite Loss
The RBT’s dual predictive capacity is its key advantage, enabling precise reward redistribution to individual instances and forecasting the next state. This leverages environmental dynamics for enhanced reward distribution as experimentally evidenced in Section 5. Integrated with off-the-shelf RL algorithms such as SAC Haarnoja et al. (2018), the RBT can enhance policy learning within the BRMDP framework, as outlined in Algorithm 1.
@@ 5 Experiment
In the following experiment section, we scrutinize the efficacy of our proposed method using benchmark tasks from both the MuJoCo Brockman et al. (2016) and the DeepMind Control Suite Tassa et al. (2018) environments, focusing on scenarios with bagged rewards. Initially, we assess the performance of our method to understand its overall effectiveness. Subsequently, we examine whether the proposed RBT reward model accurately predicts rewards. Finally, we evaluate the indispensability of each component of the reward model, questioning if every part is essential.
5.1 Performance Comparison
5.1.1 Experiment Setting
We evaluated our method on benchmark tasks from the MuJoCo locomotion suite (Ant-v2, Hopper-v2, HalfCheetah-v2, and Walker2d-v2) and the DeepMind Control Suite (cheetah-run, quadruped-walk, fish-upright, cartpole-swingup, ball_in_cup-catch, and reacher-hard). Differing from standard environments where rewards are assigned at each step, our approach involved assigning a cumulative reward at the end of each bag while assigning a reward of zero to all other state-action pairs within the bag. The maximum length for each episode was fixed at 1000 steps across all tasks.
Figure 3: Performance comparison across fixed-length reward bag in MuJoCo (top row) and DeepMind Control Suite (bottom row) environments with six different length bag settings (5, 25, 50, 100, 200, and 500) and trajectory feedback (labeled as $9$). The mean and standard deviation are displayed over 6 trials with different random seeds across a total of 1e6 time steps.
5.1.2 Baselines
In the comparative analysis, our framework was rigorously evaluated against several leading algorithms in the domain of RL with delayed reward:
@@ 5.1.2 Baselines
  • IRCR Gangwani et al. (2020): It adopted a non-parametric uniform reward redistribution approach. We have adapted IRCR for bagged reward setting.
@@
  • HC Han et al. (2022): The HC-decomposition framework was utilized to train the policy using a value function that operates on sequences of data. We employed the code as provided by the original paper.
  • Shaping Hu et al. (2020): A widely used reward shaping method for sparse reward setting, the shaping reward function is adopted from Hu et al. (2020).
@@
While methods like RUDDER Arjona-Medina et al. (2019) and Align-RUDDER Patil et al. (2022) are known for addressing the problem of trajectory feedback, previous studies Gangwani et al. (2020); Ren et al. (2021); Zhang et al. (2023) have shown superior performance using referenced methods. Additionally, since Align-RUDDER relies on successful trajectories for scoring state-action pairs, which is impractical in MuJoCo Patil et al. (2022), we ultimately excluded both methods from our comparison. Sparse reward baselines were included in our comparisons to highlight the distinctions and challenges unique to our bagged reward setting. While both settings provide rewards after a sequence, sparse rewards are Markovian, whereas our bagged rewards are non-Markovian. Therefore, including sparse reward baselines illustrates that directly applying sparse reward methods to bagged rewards is possible but not effective. For more discussion on sparse rewards, see “Discussion on Sparse Rewards” in Appendix. Besides, detailed descriptions of the model parameters and hyper-parameters used during training are provided in “Experiment Settings and Implementation Details”, with more experimental results available in “Additional Experimental Results” in Appendix.
5.1.3 Evaluation Metric
We report the average accumulative reward across 6 seeds with random initialization to demonstrate the performance of evaluated methods. Higher accumulative reward in evaluation indicates better performance.
5.1.4 Experimental Results
@@
In the fixed-length reward bag experiment (see Fig. 3), we conducted experiments with six bag lengths (5, 25, 50, 100, 200, and 500) and trajectory feedback (labeled as $9$) across each environment, where bagged reward is made by accumulating immediate rewards within bags. This aimed to illustrate the influence of varying bag lengths, providing insight into how bag size affected the performance of the learning algorithm. The SAC method, using bagged rewards directly from the environment, suffers from a lack of guidance in agent training due to missing reward information. This issue worsens with longer bag lengths, indicating that increased reward sparsity leads to less effective policy optimization. The IRCR and RRD methods, treating rewards uniformly within a reward bag, outperform SAC, suggesting benefits from even approximate guidance. However, notable variance in their results indicates potential consistency and reliability issues. The HC method excels only with shorter bag lengths, suggesting that this value function modification method struggles to utilize information from longer sequences. The Shaping and LIRPG methods exhibit subdued performance across tasks, as they are designed to solve sparse reward problems with Markovian rewards, which do not align with the reward bag setting. The proposed RBT method consistently outperforms the other approaches across all the environments and bag lengths, showing that it is not only well-suited for environments with short reward bags but also capable of handling large reward bag scenarios. This demonstrated the capability of RBT to learn from the sequence of instances and, by integrating bagged reward information, accurately allocate rewards to instances, thereby guiding better policy training.
@@
To validate the effectiveness of our approach under more complex conditions, we designed an experiment with arbitrary-length reward bags, allowing for overlaps or gaps between reward bags. This setup simulated more realistic scenarios and tested the robustness of our method. The results, detailed in Table 1, confirm the superior performance of the proposed RBT method in these complex reward settings. These findings suggest RBT’s potential for broader applications, demonstrating its versatility and robustness in handling intricate reward dynamics.
We also conducted experiments in sparse reward environments, where tasks are rewarded with 1 at the end of the trajectory if completed and 0 otherwise. In this setting, the reward does not accumulate as in our redistribution model. As shown in Table 2, RLTF methods do not perform well in sparse reward environments. Our proposed method not only outperforms RLTF baselines but also exceeds the performance of sparse reward baselines.
Table 2: Performance comparison on sparse reward environments over 6 trials with 1e6 time steps for training, presenting average scores and standard deviations. Rewards are given with 1 at the end of the trajectory if completed and 0 otherwise. The best and comparable methods based on the paired t-test at the significance level $5\%$ are highlighted in boldface.
Environment SAC IRCR RRD HC Shaping LIRPG RBT(ours)
ball_in_cup-catch 646.90 773.19 666.18 76.19 847.88 125.50 972.19
(231.15) (73.06) (33.02) (67.25) (178.01) (21.25) (5.50)
reacher-hard 531.88 584.76 347.16 7.57 664.49 8.74 735.67
(127.94) (269.32) (106.56) (3.50) (170.50) (2.36) (144.90)
@@ 6 Conclusion
In this paper, we introduced a novel learning framework called Reinforcement Learning from Bagged Rewards (RLBR). To address the challenges posed by this framework, we established theoretical connections between Bagged Reward MDPs (BRMDPs) and MDPs after reward redistribution, providing a solid foundation for our approach. Based on this theory, we proposed a reward model, the Reward Bag Transformer (RBT), designed to efficiently redistribute rewards by interpreting contextual information within bags and understanding environmental dynamics. The efficacy of RBT was demonstrated through extensive experiments, where it consistently outperformed existing methods across various reward bag scenarios. Additionally, our case studies highlighted RBT’s ability to effectively reallocate rewards while maintaining fidelity to the underlying true reward structure.
@@ Appendix A Broader Impact
In this work, we introduce the problem of Reinforcement Learning from Bagged Reward (RLBR), and propose a Transformer-based reward model that uses bidirectional attention to interpret context within a bag and accurately allocate rewards to each instance. On the one hand, we recognize that these techniques could raise some potential issues. As collecting bagged rewards is much more convenient and natural than gathering instance-level rewards, this could lead to some risks of abusing unauthorized data. On the other hand, we believe that developing these techniques is still necessary for the importance of solving reinforcement learning tasks with bag-level feedback. Furthermore, there are many techniques for preserving data privacy, which can be compatible with our approach to avoid such problems.